OpenStudy (anonymous):

x'' + 4x = 0 ; x(0) = 5, x'(0) = 0 . use laplace to solve ivp

hartnn (hartnn):

did u try this ? what u got ?

OpenStudy (anonymous):

i got something with -5s and 4/s^2 and its 4cos2t or something

OpenStudy (anonymous):

answer is 5cos2t haha but im having a hard time undertstanding the book

hartnn (hartnn):

L[x''] = ... ?

OpenStudy (anonymous):

so L(x'' + 4x) = L(0) ?

OpenStudy (anonymous):

if i can just figure out how to do this one then i can figure out the rest

hartnn (hartnn):

ok, L[x''] = s^2 F(x)-s x(0)-x'(0)

OpenStudy (anonymous):

how do we know this? and is it s^2*F(x) - s *x() - x'(0) ?

OpenStudy (anonymous):

and what is F(x) haha

hartnn (hartnn):

thats general formula and F(x) .... i had to write X(s) is the L.T of x.

OpenStudy (anonymous):

okay so its L[x''] = s^2*X(s) - s*x(0) - x'(0) ?

hartnn (hartnn):

yup, and you are given x'(0)

hartnn (hartnn):

and x(0)

OpenStudy (anonymous):

im so lost still

OpenStudy (anonymous):

dont you have to find the L of 4x too?

hartnn (hartnn):

yes, ofcourse.

hartnn (hartnn):

it would be 4 X(s)

hartnn (hartnn):

x'' + 4x = 0 taking L.T s^2*X(s) - s*x(0) - x'(0) + 4 X(s) =0

OpenStudy (anonymous):

okay so L = s^2 *X(s of x) - sx(0) - x'(0) + 4X(s of x) ?

hartnn (hartnn):

yeah, now put the given initial conditions.

OpenStudy (anonymous):

what is X(s) for the 4X(s) ? is it still x?

OpenStudy (anonymous):

how do they get 5cos2t haha

OpenStudy (anonymous):

the Lt of x is 1/s^2, right?

hartnn (hartnn):

isolate X(s) from that equation

hartnn (hartnn):

then take inverse L.T

OpenStudy (anonymous):

okay right now i have s^2X(s) - sx(0) - x'(0) + 4X(s) = 0

hartnn (hartnn):

put the initial values.

OpenStudy (anonymous):

okay so s^2X(s) - s(5) - 0 + 4X(s) = 0

hartnn (hartnn):

yes, now isolate X(s)

OpenStudy (anonymous):

s^2X(s) + 4X(s) = 5s

OpenStudy (anonymous):

= 5cos2t omg i got it! well i dont understand it but i got an answer by struggling my way through it , thank you for all your help :)

hartnn (hartnn):

what did u not understand ?

OpenStudy (anonymous):

the very first equation, where you got that

OpenStudy (anonymous):

and how to apply it

hartnn (hartnn):

thats a standard equation...like a formula.

OpenStudy (anonymous):

oh okay that makes sense

hartnn (hartnn):

L[x'] = sX(s) - x(0)

OpenStudy (anonymous):

okay and what is X(s) again?

OpenStudy (anonymous):

oh its the laplace, right?

hartnn (hartnn):

X(s) is the L.T of x.

hartnn (hartnn):

yes.

hartnn (hartnn):

L[x'] = sX(s) - x(0) L[x''] = s^2X(s) - s x(0)- x'(0) L[x'''] = s^3X(s) -s^2 x(0) - sx'(0) -x''(0) and so on...

OpenStudy (anonymous):

OH thank you so much!!!!!

hartnn (hartnn):

welcome ^_^